OpenStudy (anonymous):

how to get f^-1(3) if f(x) = x^7 + x +1 (The answer is 1, but I don't know how to get to there).

OpenStudy (anonymous):

First suppose that: \(f^{-1}(3) = x\). This means \(f(f^{-1}(3)) = f(x) \implies f(x) = 3\). So what we have is: \[ f(x) = 3 \implies 3 = x^7 + x +1 \]

OpenStudy (anonymous):

Subtracting \(3\) from both sides gives us \[ 0 = x^7+x-2 \]So you have to find the roots now.

hero (hero):

You can find it relatively easy using TI-Nspire CAS calc. I wouldn't recommend trying to find it algebraically though.

OpenStudy (anonymous):

Thank you!